OpenStudy (anonymous):

The sum of the ages of a woman and her son is 36 years. Six years from now the woman will be twice as old as her son. How old is she now?

OpenStudy (nikato):

Let the age of the woman=w And the son=s

OpenStudy (nikato):

s+w=36

OpenStudy (nikato):

Other equation 6+w= 2(s+6)

OpenStudy (nikato):

Subtract w on both sides in equation 1 So s=36-w And substitute this into second equation and solve for w